I don't understand how to solve this problem, please help me find the steps to solve this problem.
The measure of the second angle of a triangle is 50 more than that of the first. The measure of the third angle is 10 less than twice the first. Find the measure of the angles.

Answer by mananth(16949)

let one angle be x deg
second angle = x+50 deg
Third angle = 2x-10
sum of all angles of triangle = 180 deg
x+x+50+2x-10=180
4x+40=180
4x=180-40
4x= 140
/4
x=35
Plug x in the angles
first angle 35
second = 85
third = 60